Science Report – March 10th

Research report 10 March 2020

Crew 223

Submitted by Crew Scientist Marion Lebrun

1. LOAC and Weather Station:

Nothing to report. Everything is working as it should.

2. MegaARES:

Nothing to report. Everything is working as it should. We are waiting a bit longer to collect another round of data.

3. Solar oven:

The oven was taken outside a second time on sol 8. As the sun was not very bright, it only heated the 600mL of water from 12°C to 45°C during the EVA time of the day. Although the water was not as warm as the first time (due to the lack of sun light), Florian is satisfied with the collected data and the photos he took of the experiment.

4. BackPack and LCVG:

Nothing to report.

5. Foreign languages (Human Factors):

The experiment is now over (it ended on sol 7). I have noticed remarkable progress in communication within the pairs. All three teams managed to do their best performances on the last day although the LEGO model to build was the most complicated of all seven. When we come back to France, the videos and pictures taken will be sent to the Human Factors department of our school to be thoroughly analyzed.

6. Group collaboration (Human Factors):

The experiment started on sol 8. Two groups of 3 crew members play “Keep Talking and Nobody Explodes”: they have 5 minutes to defuse a bomb placed on a Martian rover. Two experts have the bomb defusing manual and need to communicate the information to the defuser who only sees the bomb screen. Just like the first experiment, this one aims at studying the communication within the teams, between the defuser and the experts. It is too early to collect relevant results.

7. Crew members’ individual planning display:

The display of the sol schedules and the hyperlink to access the available Internet data are now ready but we still cannot find a way to display a hyperlink giving direct access to Mission Support emails. This is the last thing Blandine and I would like to achieve by the end of the mission.

8. Relaxation and positive psychology (Human Factor):

Everything works well. The whole crew has been getting used to our daily relaxation time, it truly regulates our mood and gives us a bowl of positive energy to finish the day.

9. Music for plants (GreenHab):

Nothing to report. It is still too early to notice remarkable results.

10. Spirulina growth (GreenHab):

Nothing to report. Everything is working as it should, nothing new.

11. Aerospring (GreenHab):

Nothing to report.

12. Water consumption monitoring:

Aurelien realized what was wrong with the monitoring: it is the way the buttons that classify the different water usages are coded that needs to be changed. Other than that, everything else is working.

13. Supernova research:

No new observation due to bad weather but Florian still hopes the weather will improve in the next few days.

Copyright © The Mars Society. All rights reserved. | Main Site